# Fractionally modulated discrete Carleson’s Theorem and pointwise Ergodic Theorems along certain curves

### v1

@@ 1. Introduction
Lilian Pierce conjectured that the discrete analogue of $\mathcal{C}_{2}$ is bounded on $\ell^{2}({\mathbb{Z}}).$ The first result in this direction was by Krause and Lacey [15] where the $\ell^{2}({\mathbb{Z}})$-boundedness was verified under the restriction that the set of modulations has sufficiently small arithmetic Minkowski dimension. Subsequently, this restriction was dropped and the result was generalized in higher dimensions [16, 14] with the $\ell^{p}(\mathbb{Z})$ case settled as well, for $p\in(1,\infty)$.
@@ 1.3. Notation
If $A,B$ are two non-negative quantities, we write $A\lesssim B$ or $B\gtrsim A$ to denote that there exists a positive constant $C$ possibly depending on $c$ such that $A\leq CB$. We write $A\simeq B$ to denote that $A\lesssim B$ and $A\gtrsim B$. In the following sections we write $e(x)$ instead of $e^{2\pi ix}$. We denote the torus $\mathbb{R}/\mathbb{Z}$ by $\mathbb{T}$ and as usual we identify it with $[-1/2,1/2)$, so that $\|\xi\|=|\xi|$, where $\|\xi\|=\min\{|\xi-n|:\,n\in\mathbb{Z}\}$. Finally, for any $d\in\mathbb{N}$ and any finitely supported function $f\colon\mathbb{Z}^{d}\to\mathbb{C}$ we define the Fourier transform of $f$ by
@@
$\mathcal{F}_{\mathbb{R}^{d}}[f](\xi)=\int_{\mathbb{R}^{d}}f(x)e(-\xi\cdot x){\rm d}x\quad\text{and}\quad\mathcal{F}^{-1}_{\mathbb{R}^{d}}[f](x)=\int_{\mathbb{R}^{d}}f(\xi)e(\xi\cdot x){\rm d}\xi\quad\text{ respectively.}$.
@@
$[t]_{\mathsf{i}}^{c}=\mathsf{sign}(t)^{\mathsf{i}}|t|^{c}=\left\{\begin{array}[]{ll}|t|^{c}\text{,}&\text{if }\mathsf{i}=0\\ \mathsf{sign}(t)|t|^{c}\text{,}&\text{if }\mathsf{i}=1\text{.}\\ \end{array}\right.$
@@ Acknowledgment
We would like to thank Michael Lacey and Mariusz Mirek for their constant support and helpful discussions.
@@ 2. Kernel Decomposition and Exponential Sum Estimates
$\sum_{n\neq 0}f(x-n)\frac{e(\lambda\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor)}{n}=\int_{\mathbb{T}}\widehat{f}(\xi)e(\xi x)\left(\sum_{n\neq 0}\frac{e(\lambda\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor-\xi n)}{n}\right){\rm d}\xi$
$=\sum_{j\in\mathbb{N}}\int_{\mathbb{T}}\widehat{f}(\xi)e(\xi x)\left(\sum_{n\neq 0}e(\lambda\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor-\xi n)\psi_{j}(n)\right){\rm d}\xi=\sum_{j\in\mathbb{N}}\int_{\mathbb{T}}\widehat{f}(\xi)e(\xi x)m_{j}^{\mathsf{i}}(\xi,\lambda){\rm d}\xi\text{.}$
@@
For every $c\in(1,2)$ there exists a positive constant $C_{c}$ such that for all $N,M\in\mathbb{N}$ and $\xi_{1},\xi_{2}\in[-1/2,1/2)\setminus\{0\}$ we have
$\Big{|}\sum_{n=1}^{N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\Big{|}\leq C_{c}\min\big{\{}\big{(}1+\log(M)\big{)}\big{(}NM^{-1}+N^{\frac{c}{2}}M^{\frac{1}{2}}+N^{1-\frac{c}{2}}|\xi_{2}|^{-\frac{1}{2}}\big{)},|\xi_{1}|^{-1}(N^{c}|\xi_{2}|+1)\big{\}}\text{.}$
@@
We begin with the first bound. We estimate dyadic parts of the sum: let $P,P^{\prime}\in\mathbb{N}$ be such that $P^{\prime}\in[P,2P]$, then we have that
@@
We have that for any $M\in\mathbb{N}$, $20<|x|\leq 1/$ 2 and $y\in\mathbb{R}$
@@
$g_{M,\xi_{2}}(n)=e(-\xi_{2}\{n^{c}\})-\sum_{|m|\leq M}c_{m}(\xi_{2})e(mn^{c})=O\bigg{(}\min\bigg{\{}1,\frac{1}{M\|n^{c}\|}\bigg{\}}\bigg{)}$
@@
$\bigg{|}\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\bigg{|}$
(2.3) $\leq\bigg{|}\sum_{P\leq n\leq P^{\prime}}g_{M,\xi_{2}}(n)e(\xi_{2}n^{c}+\xi_{1}n)\bigg{|}+\bigg{|}\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}n^{c}+\xi_{1}n)\sum_{|m|\leq M}c_{m}(\xi_{2})e(mn^{c})\bigg{|}\text{.}$
@@
[…]
We may use Lemma 4.3 and Lemma 4.4 from [10] to obtain that for $t\neq 0$ and $\xi\in[-1/2,1/2)$ we have
$U_{P,P^{\prime}}(t,\xi)\lesssim P^{\frac{c}{2}}|t|^{\frac{1}{2}}+P^{1-\frac{c}{2}}|t|^{-\frac{1}{2}}\quad\text{and}\quad V_{P,P^{\prime}}(M)\lesssim(1+\log M)(M^{-1}P+P^{\frac{c}{2}}M^{\frac{1}{2}})\text{.}$
For the first summand of (2.3) we immediately see that
$\bigg{|}\sum_{P\leq n\leq P^{\prime}}g_{M,\xi_{2}}(n)e(\xi_{2}n^{c}+\xi_{1}n)\bigg{|}\lesssim V_{P,P^{\prime}}(M)\lesssim(1+\log M)(M^{-1}P+P^{\frac{c}{2}}M^{\frac{1}{2}})\text{,}$
@@
(2.4) $\bigg{|}\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}n^{c}+\xi_{1}n)\sum_{|m|\leq M}c_{m}(\xi_{2})e(mn^{c})\bigg{|}=\bigg{|}\sum_{|m|\leq M}c_{m}(\xi_{2})\sum_{P\leq n\leq P^{\prime}}e((\xi_{2}+m)n^{c}+\xi_{1}n)\bigg{|}\\ \leq\sum_{|m|\leq M}|c_{m}(\xi_{2})|\Big{|}\sum_{P\leq n\leq P^{\prime}}e((\xi_{2}+m)n^{c}+\xi_{1}n)\Big{|}=\sum_{|m|\leq M}|c_{m}(\xi_{2})|U_{P,P^{\prime}}(\xi_{2}+m,\xi_{1})\\ \leq\bigg{(}\sum_{|m|\leq M}|c_{m}(\xi_{2})|\bigg{)}\sup_{|\xi_{2}|\leq|t|\leq M+1/2}U_{P,P^{\prime}}(t,\xi_{1})\lesssim(1+\log M)\sup_{|\xi_{2}|\leq|t|\leq M+1/2}\Big{(}P^{\frac{c}{2}}|t|^{\frac{1}{2}}+P^{1-\frac{c}{2}}|t|^{-\frac{1}{2}}\Big{)}\\ \lesssim(1+\log M)\big{(}P^{\frac{c}{2}}M^{\frac{1}{2}}+P^{1-\frac{c}{2}}|\xi_{2}|^{-\frac{1}{2}}\Big{)}\text{.}$
@@
$\bigg{|}\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\bigg{|}\lesssim(1+\log M)\big{(}M^{-1}P+P^{\frac{c}{2}}M^{\frac{1}{2}}+P^{1-\frac{c}{2}}|\xi_{2}|^{-\frac{1}{2}}\Big{)}\text{.}$
@@
(2.5) $\bigg{|}\sum_{1\leq n\leq N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\bigg{|}\lesssim\sum_{l=0}^{\lfloor\log_{2}N\rfloor}\Big{|}\sum_{2^{l}\leq n\leq\min(2^{l+1}-1,N)}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\Big{|}\\ \lesssim\sum_{l=0}^{\log_{2}N}(1+\log M)(M^{-1}2^{l}+2^{lc/2}M^{1/2}+2^{l(1-c/2)}|\xi_{2}|^{-1/2})\\ \lesssim(1+\log M)\big{(}NM^{-1}+N^{\frac{c}{2}}M^{1/2}+N^{1-\frac{c}{2}}|\xi_{2}|^{-1/2}\big{)}\text{,}$
and the proof of the first estimate is complete.
@@
(2.6) $\Big{|}\sum_{n=1}^{N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\Big{|}\leq\Big{|}\sum_{n=1}^{N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)-e(\xi_{1}n)\Big{|}+\Big{|}\sum_{n=1}^{N}e(\xi_{1}n)\Big{|}\\ \lesssim\Big{|}\sum_{n=1}^{N}e(\xi_{1}n)\big{(}e(\xi_{2}\lfloor n^{c}\rfloor)-1\big{)}\Big{|}+|\xi_{1}|^{-1}\\ \lesssim\Big{|}\sum_{n=1}^{N}e(\xi_{1}n)\Big{|}\cdot|\big{(}e(\xi_{2}\lfloor N^{c}\rfloor)-1\big{)}|+\Big{|}\sum_{n=1}^{N-1}\sum_{k=1}^{n}e(\xi_{1}k)\big{(}e(\xi_{2}\lfloor n^{c}\rfloor)-e(\xi_{2}\lfloor(n+1)^{c}\rfloor)\big{)}\Big{|}+|\xi_{1}|^{-1}\\ \lesssim|\xi_{1}|^{-1}+\sum_{n=1}^{N}|\xi_{1}|^{-1}|\xi_{2}|\big{(}(n+1)^{c}-n^{c}\big{)}\lesssim|\xi_{1}|^{-1}\big{(}1+|\xi_{2}|N^{c}\big{)}\text{,}\quad\text{as desired.}$
As an immediate corollary we obtain the following estimate which can be verified after a moment’s reflection by combining Proposition 2.1 and Lemma 5.5 from [6].
@@
For every $c\in(1,2)$ and $\mathsf{i}\in\left\{0,1\right\}$ there exists a positive constant $C_{c}$ such that for all $j\in\mathbb{N}$ and $\xi_{1},\xi_{2}\in[-1/2,1/2)\setminus\{0\}$ we have
$\Big{|}\sum_{n\in{\mathbb{Z}}}e(\xi_{2}\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor+\xi_{1}n)\psi_{j}(n)\Big{|}\leq C_{c}2^{-j}\min\big{\{}j\big{(}2^{(\frac{c+1}{3})j}+2^{(1-\frac{c}{2})j}|\xi_{2}|^{-\frac{1}{2}}\big{)},|\xi_{1}|^{-1}(2^{cj}|\xi_{2}|+1)\big{\}}\text{.}$
It is clear that in both propositions above we do not have to exclude the cases where only one frequency is 00. More specifically, if $\xi_{2}=0$ and $\xi_{1}\neq 0$, then the second estimate remains true. Similarly, for $\xi_{2}\neq 0$ and $\xi_{1}=0$, the first estimate holds.

### v2

@@ 1. Introduction
Lillian Pierce conjectured that the discrete analogue of $\mathcal{C}_{2}$ is bounded on $\ell^{2}({\mathbb{Z}}).$ The first result in this direction was by Krause and Lacey [15] where the $\ell^{2}({\mathbb{Z}})$-boundedness was verified under the restriction that the set of modulations has sufficiently small arithmetic Minkowski dimension. Subsequently, this restriction was dropped, the result was generalized in higher dimensions and in the multiparameter setting [16, 17, 14] with the $\ell^{p}$ case settled as well, for $p\in(1,\infty)$.
@@
1.3. Short remarks on the nonuniformity of the constants
Motivated by the atypically large, in some sense optimal for such arguments, range of exponents $c$ for which Theorem 1 is established, we wish to make some brief comments on the constant $C_{c,p}$ in the estimates (1.2). Theorem 1 holds for $c\in[1,2]$, with the case $c=1$ corresponding essentially to the $L^{p}$-boundedness of the Carleson operator and the case $c=2$ to the $\ell^{2}$-boundness of the discrete quadratic Carleson operator.
It will be clear that the constant $C_{c,p}$ for Theorem 1 guaranteed by our proof blows up as $c\to 1^{+}$ or $c\to 2^{-}$. At the expense of making our estimates slightly more cumbersome, we have quantified this nonuniformity for p=2p=2111As discussed in the previous subsection the constant $C_{c,p}$ for $p\neq 2$ is obtained through various interpolative arguments with $\ell^{2}$-estimates, and the blowup will propagate accordingly. For the sake of keeping the exposition more reasonable, and since we believe the blowup (1.3) is far from optimal, we have confined ourselves in quantifying the failure of nonuniformity with respect to $c\in(1,2)$ only for the case $p=2$., and the more precise estimate our argument affords us amounts to
(1.3) $C_{c,2}\lesssim(c-1)^{-\frac{1}{2}}(2-c)^{-\frac{13}{4}}\text{.}$
where the implied constant is absolute. Both blowups are imposed by the method essentially due to applications of van den Corput lemma, either for trigonometric sums or oscillatory integrals.
This is not suprising since the proposed method here exploits the range $c\in(1,2)$ in two different ways. On the hand, the so-called lack of arithmetical obstructions for the exponential sums with phases $(\xi_{2}\lfloor|n|^{c}\rfloor+\xi_{1}n)_{n\in\{1,\dotsc,N\}}$ for $c\in(1,2)$ allows us to compare the corresponding operators with their continuous counterpart without additional number-theoretic considerations, namely the appearence of certain multi-frequency phenomena and Gauss sums, present for $c=2$. On the other hand, since $c>1$, a certain kind of curvature is introduced, and modulation-invariance phenomena which ought to be addressed for the case $c=1$, yielding the Carleson operator, are avoided.
A rather interesting natural question is whether the constant of Theorem 1 can be chosen independently of $c\in[1,2]$. An analysis affording uniformity in the estimates of such strength most likely should not only take into account the way in which arithmetical obstructions appear as $c\to 2^{-}$ in a very quantitative manner, but also establish the continuous analogue of the result, namely, the main result from [8], with uniform constants for $c\in[1,2]$, an endeavor closely realated to Carleson’s theorem, see also Question 1.4 in [8]. Both tasks seem to go beyond the current technology and certainly beyond the methods of the present work.
@@ 1.4. Notation
If $A,B$ are two non-negative quantities, we write $A\lesssim B$ or $B\gtrsim A$ to denote that there exists an absolute positive constant $C$ such that $A\leq CB$. We write $A\simeq B$ to denote that $A\lesssim B$ and $A\gtrsim B$. In the following sections we write $e(x)$ instead of $e^{2\pi ix}$. For every $r\in\mathbb{R}$ we let $\mathbb{N}_{\geq r}\coloneqq\mathbb{N}\cap[r,\infty)$. We denote the torus $\mathbb{R}/\mathbb{Z}$ by $\mathbb{T}$ and as usual we identify it with [1/2,1/2)[-1/2,1/2), so that $\|\xi\|=|\xi|$, where $\|\xi\|=\min\{|\xi-n|:\,n\in\mathbb{Z}\}$. Finally, for any $d\in\mathbb{N}$ and any finitely supported function $f\colon\mathbb{Z}^{d}\to\mathbb{C}$ we define the Fourier transform of $f$ by
@@
$\mathcal{F}_{\mathbb{R}^{d}}[f](\xi)=\int_{\mathbb{R}^{d}}f(x)e(-\xi\cdot x){\rm d}x\quad\text{and}\quad\mathcal{F}^{-1}_{\mathbb{R}^{d}}[f](x)=\int_{\mathbb{R}^{d}}f(\xi)e(\xi\cdot x){\rm d}\xi\quad\text{ respectively.}$
@@
$t]_{\mathsf{i}}^{c}=\mathsf{sign}(t)^{\mathsf{i}}|t|^{c}=\left\{\begin{array}[]{ll}|t|^{c}\text{,}&\text{if }\mathsf{i}=0\\ \mathsf{sign}(t)|t|^{c}\text{,}&\text{if }\mathsf{i}=1\text{.}\\ \end{array}\right$.
@@ Acknowledgment
We would like to thank Michael Lacey and Mariusz Mirek for their constant support and helpful discussions. The authors are also grateful to the referee for their careful reading of the manuscript and useful remarks that led to the improvement of the presentation.
@@ 2. Kernel Decomposition and Exponential Sum Estimates
(2.2) $\sum_{n\neq 0}f(x-n)\frac{e(\lambda\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor)}{n}=\int_{\mathbb{T}}\widehat{f}(\xi)e(\xi x)\left(\sum_{n\neq 0}\frac{e(\lambda\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor-\xi n)}{n}\right){\rm d}\xi\\ =\sum_{j\in\mathbb{N}}\int_{\mathbb{T}}\widehat{f}(\xi)e(\xi x)\left(\sum_{n\neq 0}e(\lambda\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor-\xi n)\psi_{j}(n)\right){\rm d}\xi=\sum_{j\in\mathbb{N}}\int_{\mathbb{T}}\widehat{f}(\xi)e(\xi x)m_{j}^{\mathsf{i}}(\xi,\lambda){\rm d}\xi\text{.}$
@@
There exists an absolute positive constant $C$ such that for every $c\in(1,2)$ and all $N,M\in\mathbb{N}$ and $\xi_{1},\xi_{2}\in[-1/2,1/2)\setminus\{0\}$ we have
(2.3) $\Big|\sum_{n=1}^{N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\Big|\leq C(2-c)^{-1}(c-1)^{-\frac{1}{2}}\cdot\\ \cdot\min\big\{\big(1+\log(M)\big)\big(NM^{-1}+N^{\frac{c}{2}}M^{\frac{1}{2}}+N^{1-\frac{c}{2}}|\xi_{2}|^{-\frac{1}{2}}\big),|\xi_{1}|^{-1}(N^{c}|\xi_{2}|+1)\big\}\text{.}$
@@
We begin with the first bound. We estimate dyadic parts of the sum: let $P,P^{\prime}\in\mathbb{N}_{\geq 2}$ be such that $P^{\prime}\in[P,2P]$, then
@@
We have that for any $M\in\mathbb{N}$, $0<|x|\leq 1/2$ and $y\in\mathbb{R}$
@@
$g_{M,\xi_{2}}(n)=e(-\xi_{2}\{n^{c}\})-\sum_{|m|\leq M}c_{m}(\xi_{2})e(mn^{c})=O\bigg(\min\bigg\{1,\frac{1}{M\|n^{c}\|}\bigg\}\bigg)\text{,}$
@@
(2.5) $\bigg|\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\bigg|\\ \leq\bigg|\sum_{P\leq n\leq P^{\prime}}g_{M,\xi_{2}}(n)e(\xi_{2}n^{c}+\xi_{1}n)\bigg|+\bigg|\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}n^{c}+\xi_{1}n)\sum_{|m|\leq M}c_{m}(\xi_{2})e(mn^{c})\bigg|\text{.}$
@@
$V_{P,P^{\prime}}(M)=\sum_{P\leq n\leq P^{\prime}}\min\Big\{1,\frac{1}{M\|n^{c}\|}\Big\}\text{,}\quad\text{for $M\in\mathbb{N}$ }\text{.}$
Applying van der Corput lemma, see for example Corollary 8.13 in [11], yields that
(2.6) $U_{P,P^{\prime}}(t,\xi)\lesssim P^{\frac{c}{2}}|t|^{\frac{1}{2}}+(c-1)^{-\frac{1}{2}}P^{1-\frac{c}{2}}|t|^{-\frac{1}{2}}\text{,}$
which we may use in an identical manner as in the proof of Lemma 4.4 in [10] to obtain
(2.7) $V_{P,P^{\prime}}(M)\lesssim(c-1)^{-\frac{1}{2}}(1+\log M)\big(M^{-1}P+P^{\frac{c}{2}}M^{\frac{1}{2}}\big)\text{.}$
For the first summand in the second line of (2.5) we immediately see that
$\bigg|\sum_{P\leq n\leq P^{\prime}}g_{M,\xi_{2}}(n)e(\xi_{2}n^{c}+\xi_{1}n)\bigg|\lesssim V_{P,P^{\prime}}(M)\lesssim(c-1)^{-\frac{1}{2}}(1+\log M)(M^{-1}P+P^{\frac{c}{2}}M^{\frac{1}{2}})\text{,}$
@@
(2.8) $\bigg|\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}n^{c}+\xi_{1}n)\sum_{|m|\leq M}c_{m}(\xi_{2})e(mn^{c})\bigg|=\bigg|\sum_{|m|\leq M}c_{m}(\xi_{2})\sum_{P\leq n\leq P^{\prime}}e((\xi_{2}+m)n^{c}+\xi_{1}n)\bigg|\\ \leq\sum_{|m|\leq M}|c_{m}(\xi_{2})|U_{P,P^{\prime}}(\xi_{2}+m,\xi_{1})\leq\bigg(\sum_{|m|\leq M}|c_{m}(\xi_{2})|\bigg)\sup_{|\xi_{2}|\leq|t|\leq M+1/2}U_{P,P^{\prime}}(t,\xi_{1})\\ \lesssim(1+\log M)\sup_{|\xi_{2}|\leq|t|\leq M+1/2}\Big(P^{\frac{c}{2}}|t|^{\frac{1}{2}}+(c-1)^{-\frac{1}{2}}P^{1-\frac{c}{2}}|t|^{-\frac{1}{2}}\Big)\lesssim(c-1)^{-\frac{1}{2}}(1+\log M)\Big(P^{\frac{c}{2}}M^{\frac{1}{2}}+P^{1-\frac{c}{2}}|\xi_{2}|^{-\frac{1}{2}}\Big)\text{.}$
@@
(2.9) $\bigg|\sum_{P\leq n\leq P^{\prime}}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\bigg|\lesssim(c-1)^{-\frac{1}{2}}(1+\log M)\Big(M^{-1}P+P^{\frac{c}{2}}M^{\frac{1}{2}}+P^{1-\frac{c}{2}}|\xi_{2}|^{-\frac{1}{2}}\Big)\text{.}$
@@
(2.10) $\bigg|\sum_{1\leq n\leq N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\bigg|\lesssim 1+\sum_{l=1}^{\lfloor\log_{2}N\rfloor}\Big|\sum_{2^{l}\leq n\leq\min(2^{l+1}-1,N)}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\Big|\\ \lesssim 1+(c-1)^{-\frac{1}{2}}\sum_{l=0}^{\lfloor\log_{2}N\rfloor}(1+\log M)(M^{-1}2^{l}+2^{lc/2}M^{1/2}+2^{l(1-c/2)}|\xi_{2}|^{-1/2})\\ \lesssim\frac{1}{(2-c)\sqrt{c-1}}\cdot(1+\log M)\big(NM^{-1}+N^{\frac{c}{2}}M^{1/2}+N^{1-\frac{c}{2}}|\xi_{2}|^{-1/2}\big)\text{,}$
where for the final estimate we used that for every $\theta>0$ we have
(2.11) $\sum_{l=0}^{\lfloor\log_{2}N\rfloor}2^{l\theta}=\frac{\big(2^{\theta}\big)^{\lfloor\log_{2}N\rfloor}-1}{2^{\theta}-1}\leq\frac{N^{\theta}}{e^{\theta\log 2}-1}\leq\frac{N^{\theta}}{\theta\log(2)}\text{,}\quad\text{since $e^{x}-1\geq x$.}$
@@
(2.12) $\Big|\sum_{n=1}^{N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)\Big|\leq\Big|\sum_{n=1}^{N}e(\xi_{2}\lfloor n^{c}\rfloor+\xi_{1}n)-e(\xi_{1}n)\Big|+\Big|\sum_{n=1}^{N}e(\xi_{1}n)\Big|\\ \lesssim\Big|\sum_{n=1}^{N}e(\xi_{1}n)\big(e(\xi_{2}\lfloor n^{c}\rfloor)-1\big)\Big|+|\xi_{1}|^{-1}\\ \lesssim\Big|\sum_{n=1}^{N}e(\xi_{1}n)\Big|\cdot|\big(e(\xi_{2}\lfloor N^{c}\rfloor)-1\big)|+\Big|\sum_{n=1}^{N-1}\sum_{k=1}^{n}e(\xi_{1}k)\big(e(\xi_{2}\lfloor n^{c}\rfloor)-e(\xi_{2}\lfloor(n+1)^{c}\rfloor)\big)\Big|+|\xi_{1}|^{-1}\\ \lesssim|\xi_{1}|^{-1}+\sum_{n=2}^{N}|\xi_{1}|^{-1}|\xi_{2}|\big((n+1)^{c}-(n-1)^{c}\big)\lesssim|\xi_{1}|^{-1}\big(1+|\xi_{2}|N^{c}\big)\text{,}\quad\text{as desired.}$
As an immediate corollary we obtain the following estimate.
@@
There exists an absolute positive constant $C$ such that for every $c\in(1,2)$, $\mathsf{i}\in\left\{0,1\right\}$ and all $j\in\mathbb{N}_{\geq(2-c)^{-1}}$ and $\xi_{1},\xi_{2}\in[-1/2,1/2)\setminus\{0\}$ we have
(2.13) $\Big|\sum_{n\in{\mathbb{Z}}}e(\xi_{2}\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor+\xi_{1}n)\psi_{j}(n)\Big|\\ \leq C2^{-j}\min\big\{(c-1)^{-\frac{1}{2}}(2-c)j\big(2^{(\frac{c+1}{3})j}+2^{(1-\frac{c}{2})j}|\xi_{2}|^{-\frac{1}{2}}\big),|\xi_{1}|^{-1}(2^{cj}|\xi_{2}|+1)\big\}\text{.}$
Proof.
Note that for every such $j\in\mathbb{N}$ and $P^{\prime}\in[2^{j-3},2^{j-1}]$ we may apply the estimate (2.9) for $M=\lfloor 2^{\frac{2-c}{3}j}\rfloor$ to obtain
(2.14) $\Big|\sum_{2^{j-3}\leq n\leq P^{\prime}}e(\xi_{2}\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor+\xi_{1}n)\Big|\lesssim(c-1)^{-\frac{1}{2}}(1+(2-c)j)\Big(\lfloor 2^{\frac{2-c}{3}j}\rfloor^{-1}2^{j}+2^{\frac{c}{2}j}2^{\frac{2-c}{6}j}+2^{(1-\frac{c}{2})j}|\xi_{2}|^{-\frac{1}{2}}\Big)\\ \lesssim(c-1)^{-\frac{1}{2}}(2-c)j\Big(2^{(\frac{c+1}{3})j}+2^{(1-\frac{c}{2})j}|\xi_{2}|^{-\frac{1}{2}}\Big)\text{,}$
where we used that $\lfloor 2^{\frac{2-c}{3}j}\rfloor\geq 2^{\frac{2-c}{3}j}-1\geq\Big(\frac{2^{\frac{1}{3}}-1}{2^{\frac{1}{3}}}\Big)2^{\frac{2-c}{3}j}$, where the last inequality is true since it is equivalent to
$\frac{1}{2^{\frac{1}{3}}}2^{\frac{2-c}{3}j}\geq 1\iff 2^{\frac{(2-c)j}{3}}\geq 2^{\frac{1}{3}}\text{,}\quad\text{which is true by our assumption.}$
Using triangle inequality, breaking the sum above as follows $\big|\sum_{2^{j-3}\leq n\leq P^{\prime}}\big|\leq\big|\sum_{1\leq n\leq P^{\prime}}\big|\leq+\big|\sum_{1\leq n<2^{j-3}}\big|$ and using the estimate (2.12) yields
$\Big|\sum_{2^{j-3}\leq n\leq P^{\prime}}e(\xi_{2}\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor+\xi_{1}n)\Big|\lesssim|\xi_{1}|^{-1}\big(2^{cj}|\xi_{2}|+1\big)\text{.}$
Combining the two estimates above with Lemma 5.5 from [6] yields the following
(2.15) $\Big|\sum_{n\geq 0}e(\xi_{2}\mathsf{sign}(n)^{\mathsf{i}}\lfloor|n|^{c}\rfloor+\xi_{1}n)\psi_{j}(n)\Big|\\ \lesssim 2^{-j}\min\big\{(c-1)^{-\frac{1}{2}}(2-c)j\big(2^{(\frac{c+1}{3})j}+2^{(1-\frac{c}{2})j}|\xi_{2}|^{-\frac{1}{2}}\big),|\xi_{1}|^{-1}(2^{cj}|\xi_{2}|+1)\big\}\text{,}$
and a symmetric argument establishes the same estimates for $n<0$. The proof is complete. ∎
It is clear that in both propositions above we do not have to exclude the cases where only one frequency is 0. More specifically, if $\xi_{2}=0$ and $\xi_{1}\neq 0$, then the second estimate remains true. Similarly, for $\xi_{2}\neq 0$ and $\xi_{1}=0$, the first estimate holds.